The point A (2,4) is joined to any point P on the circle x^2+y^2=4.The locus of the mid point of AP as p moves on circle is:

Grade:12th pass

1 Answers

Aditya Gupta
5 years ago
let p have coordinates 2cosq, 2sinq
now mid point be (h,k)
so, 2h=2+2cosq or h – 1= cosq
and 2k= 4+2sinq or k – 2=sinq
squaring and adding, nd replacing h and k by x and y resp,
(x – 1)^2+(y – 2)^2= 1 is the reqd locus